OpenStudy (anonymous):

solve for x: -5/x+1/ = 10

OpenStudy (redohawk):

is this supposed to represnt division? /

OpenStudy (anonymous):

Yeah,what is up with the lose / ???

OpenStudy (anonymous):

\(\Large \frac{-5}{x+1}=10\) \(\Large -5=10(x+1)\) \(\Large 10x+10=-5\) \(\Large 10x=-5-10\) \(\Large 10x=-15\) \(\Large x=\frac{-15}{10}=\frac{-3}{2}\)

OpenStudy (redohawk):

like a boss !!!

OpenStudy (anonymous):

its supposed to be like this Solve for x: −5|x + 1| = 10

OpenStudy (redohawk):

−5|x + 1| = 10 -5x + 1 = 10 x + 1 = 10 x = 9 is that one of the choices? x = 9

OpenStudy (redohawk):

hope i helped @lizzxoxo221

OpenStudy (anonymous):

x = 0 x = −3 and x = 1 x = −1 and x = 3 these are the choices and there is also no solution

OpenStudy (redohawk):

i a thinking no solution...

OpenStudy (redohawk):

no no no!

OpenStudy (redohawk):

x = -3 coz... -3 + 1 = -2, and... -2 * -5 = 10

OpenStudy (anonymous):

so is it c ?

OpenStudy (anonymous):

There is no solution

OpenStudy (anonymous):

it is False

OpenStudy (anonymous):

well I tried no solution and it says its not right

OpenStudy (anonymous):

If x > 0 -5x -5 = 10 which = false

OpenStudy (redohawk):

i told you x = -3 what happened?

OpenStudy (anonymous):

than is it 0 because its definitely not no solution

OpenStudy (anonymous):

wait so its b than? @RedoHawk

OpenStudy (anonymous):

If we are using real numbers \[ -5 \sqrt{(x+1)^2} = 10 \] There is no solution unless imaginary numbers but I have not worked that out yet. Let x = i and see

OpenStudy (redohawk):

sorry i i give any wrong answers :(